OpenStudy (anonymous):

Sonu invested 10% more than Mona. Mona invested 10% less than Raghu. If the total sum of their investment is Rs. 5,780, how much amount did Raghu invest ?

OpenStudy (anonymous):

\[ s = (1+0.10)m = 1.10 m \\ m = (1-0.10) r = 0.9r \\ s+m = 5780 \]

OpenStudy (anonymous):

This is a system of equations.

OpenStudy (anonymous):

thanks a lot